Note: You MUST show ALL your work to receive FULL credit- that is, show timelines and calculator sequence to justify your answers. 1. Scott Frost has just won the Cornhusker Sweepstakes. He is entitled to receive one of the following prizes: a. $100,000 per year forever with the first payment received today b. $40,000 beginning next year growing by 4% forever. c. $200,000 for ten years starting in year 6. d. Ten payments of $150,000 per year starting one year from today. e. Five payments of $300,000 per year starting one year from today, then five more payments of $300,000 beginning eleven years from today. Assuming a discount rate of 10%, evaluate all of the alternative prizes and choose the prize that would have the greatest wealth maximizing effect for Frosty. (10 points)
Answer #1

We compute the present value of these options as below

A: PV= Amount received now+ CF1/r

=100000+ 100000/10%

=1100000

B: CF1/(r-g)

= 40000/(10%-4%)

=666666.67

C: PV = (200000*P/A(10%,10))/ (1+r)^5

=200000*6.145/ (1+0.1)^5

=763112.3

D: PV= P*P/A,10%,10

= 150000*6.145

=921750

E: PV = P*P/A,10%,5 + P*P/A,10%,5 / (1+r)^10

= 300000*3.791 + 300000*3.791/1.1^10

=1575778

Option E is the best since its present value is the highest.